consider a particle system with particles having mass m1=2kg and m2=3kg in the x-y plane . if the first particle is pushed with a force 3i-4j N and the second object is pushed with force -4i+4j N, what is the magnitude and direction of the acceleration of centre of mass?

Grade:11

2 Answers

Vikas TU
6 years ago
Dear Student,
net force = 3i-4j-4i+4j = -i
       Magnitude of force = 1 unit
       Direction of force = east (along negative x axis).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Bibek
6 years ago
Friends Net force = (3-4)I+(-4+4)j=-i. that means force acts in negative x-axis.Now magnitude of total mass = 2+3=5kg.So magnitude of acceleration is equal to Net force divided by total mass i.e, a=F/m=1/5=.2m/s^2.Now direction of acceleration is same as direction of net resultant force .By Bibek , Btech 4th year ,Vssut,Burla,sambalpur,odisha